# Data Trade and Consumer Privacy

### v2

@@ A.8 Additional Results on General Service
$2\left(x_{L}-x_{H}\right)=\pi\left(l|H\right)-\pi\left(l|L\right).$ (24)
Plugging (24) into the intermediary’s objective function which maximizes
$V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\left(1-\mu_{0}\right)\Delta v$
$+\Big{[}\left(x_{H}-L\right)^{2}-\Big{(}\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}+x_{H}-L\Big{)}^{2}\Big{]}\left(1-\mu_{0}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\left(1-\mu_{0}\right)v_{H}$
subject to $OB_{H}$, $\pi\left(l|L\right)\in\left[0,1\right]$, $\pi\left(l|H\right)\in\left[0,1\right]$. The first order derivative with respect to $\pi(l|H)$ is positive, $\pi(l|H)=1$.
$x_{H}-L$ $=$ $-v_{L}-\frac{\pi(l|H)-\pi(l|L)}{2}$
$x_{H}-H$ $=$ $-\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}\left(1-\mu_{0}\right)$
from which
$x_{H}$ $=$ $\frac{H-(L-v_{L})(1-\mu_{0})}{\mu_{0}}$
$\pi(l|L)$ $=$ $1\displaystyle 1+2\times\frac{H-L+v_{L}}{\mu_{0}}>$ 1
So $\pi(l|L)=1$, and then $x_{H}=H$. By (24) $x_{L}=x_{H}=H$ which is a contradiction with $x_{L}\leq L$. Hence, it’s not feasible that only $IR_{H}$ is binding.
@@
A.9 Proof of Proposition 6
Proof.
For $\mu_{0}>\frac{v_{L}}{v_{H}}$, social welfare with data banning is $V+\mu_{0}v_{H}$, and social welfare without regulation
  1. (i)
    if $\frac{\delta^{2}}{\Delta v}\geq 1$: $V+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
  2. (ii)
  3. (iii)
each case is strictly greater than $V+\mu_{0}v_{H}$. This is obvious for the first and third cases. In the second case, $v_{L}-\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}=\frac{-(\Delta v)^{2}-\delta^{4}+2\delta^{2}(v_{L}+v_{H})}{4\delta^{2}}>0$ when $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$. Hence, banning data trade decreases social welfare for $\mu_{0}>\frac{v_{L}}{v_{H}}$.

### v3

@@ A.9 Proof of 5
which indeed satisfies $OB_{H}$ and $IC_{H}$. But the solution violates $IR_{L}$ because $2\frac{H-\left(1-\mu_{0}\right)L}{\mu_{0}}-L-H=\frac{(2-\mu_{0})(H-L)}{\mu_{0}}>0=[1-\pi(l|H)]\Delta v/\delta$. Hence, $IC_{H}$ is binding:
$2\left(x_{L}-x_{H}\right)=\left[\pi\left(l|H\right)-\pi\left(l|L\right)\right]\frac{\Delta v}{\delta}.$ (31)
Plugging (31) into the intermediary’s objective function to substitute out $x_{L}$, it becomes
$\max_{x,\pi}V-\left(x_{H}-{H}\right)^{2}$ $+\pi(l|H)(1-\beta\mu_{0})\Delta v+\pi(l|L)\beta(1-\mu_{0})v_{L}$
$+\left[\left(x_{H}-{L}\right)^{2}-\left(x_{L}-{L}\right)^{2}\right]\left(1-\mu_{0}\right)-(1-\beta\mu_{0})v_{H}-(1-\beta)v_{L}$
$=V-\left(x_{H}-{H}\right)^{2}$ $+\pi(l|H)(1-\beta\mu_{0})\Delta v+\pi(l|L)\beta(1-\mu_{0})v_{L}$
$+\left[\left(x_{H}-{L}\right)^{2}-\left(\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}\frac{\Delta v}{\delta}+x_{H}-L\right)^{2}\right]\left(1-\mu_{0}\right)$
$-(1-\beta\mu_{0})v_{H}-(1-\beta)v_{L}$
subject to $OB_{H}$, $\pi\left(l|L\right)\in\left[0,1\right]$, $\pi\left(l|H\right)\in\left[0,1\right]$. The first order derivative with respect to $\pi(l|H)$ is positive, so $OB_{H}$ is binding: $\frac{v_{L}}{v_{H}}(1-\mu_{0})\pi(l|L)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi(l|H)=\frac{v_{L}}{v_{H}}-\mu_{0}$. Let $\lambda$ be the Lagrangian multiplier of the binding $OB_{H}$ constraint.
We show that either the binding or nonbinding $\pi(l|L)\in[0,1]$ is impossible. First, suppose $\pi(l|L)\in[0,1]$ is not binding and ignore it. The first order conditions with respect to $\pi(l|L)$, $\pi(l|H)$ and $x_{H}$ are respectively
$\beta(1-\mu_{0})v_{L}+2(1-\mu_{0})\left[\frac{\pi(l|H)-\pi(l|L)}{2}\frac{\Delta v}{\delta}+x_{H}-L\right]\frac{\Delta v}{2\delta}+\lambda\frac{v_{L}}{v_{H}}(1-\mu_{0})$ $=$ $0$
$(1-\beta\mu_{0})\Delta v-2(1-\mu_{0})\left[\frac{\pi(l|H)-\pi(l|L)}{2}\frac{\Delta v}{\delta}+x_{H}-L\right]\frac{\Delta v}{2\delta}-\lambda(1-\frac{v_{L}}{v_{H}})\mu_{0}$ $=$ $0$
$-2(x_{H}-H)-(1-\mu_{0})[\pi(l|H)-\pi(l|L)]\frac{\Delta v}{\delta}$ $=$ $0$
the sum of the first two conditions yields
$\beta(1-\mu_{0})v_{L}+(1-\beta\mu_{0})\Delta v+\lambda(\frac{v_{L}}{v_{H}}-\mu_{0})=0$
which contradicts with $\mu_{0}<\frac{v_{L}}{v_{H}}$. Second, suppose $\pi(l|L)\in[0,1]$ is binding, then $\pi(l|L)=1=\pi(l|H)$ and $x_{H}=H$ by the third condition. By (31) $x_{L}=x_{H}=H$ which contradicts with $x_{L}\leq L$.
In sum, it’s not feasible that only $IR_{H}$ is binding.
@@
By plugging $f_{H}$ and $f_{L}$ into the intermediary’s objective, we can rewrite it as
$f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\beta\left[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\mu_{0}v_{H}-v_{L}\right]$
$=$ $[V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\Delta v-\Delta v]\mu_{0}+[V-\left(x_{L}-L\right)^{2}]\left(1-\mu_{0}\right)$
$+\beta\left[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\mu_{0}v_{H}-v_{L}\right]$
$=$ $V-\mu_{0}\left(x_{H}-H\right)^{2}+\mu_{0}\pi\left(l|H\right)\Delta v-\mu_{0}\Delta v-\left(1-\mu_{0}\right)\left(x_{L}-L\right)^{2}$
$+\beta\left[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\mu_{0}v_{H}-v_{L}\right]$
$=$ $V-\mu_{0}\left(x_{H}-H\right)^{2}-\left(1-\mu_{0}\right)\left(x_{L}-L\right)^{2}+\mu_{0}(1-\beta)\pi\left(l|H\right)\Delta v+\beta\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
$+\beta\mu_{0}v_{H}-\beta v_{L}-\mu_{0}\Delta v$
The intermediary maximizes the objective subject to $OB_{H}$ and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$ and
@@
Hence, given $\beta\in(0,1)$, $x_{L}=L$, $x_{H}=H$ and $\pi\left(l|L\right)=1$ and $\pi\left(l|H\right)=1$.
In sum, when $\mu_{0}<\frac{v_{L}}{v_{H}}$, $x_{L}=L$, $x_{H}=H$, $f_{L}=f_{H}=V$, $\pi\left(l|L\right)=\pi\left(l|H\right)=1$. The intermediary’s payoff is $V$; the producer’s payoff is $v_{L}$; consumers’ surplus is $\mu_{0}\Delta v$.